OpenStudy (anonymous):

Find sec(-7pi/3) without calculator.

OpenStudy (***[isuru]***):

Hi, \[\sec \theta = \frac{ 1 }{ \cos \theta }\] if that is true .. How could u write ur question in 1 over cos form ... ?

OpenStudy (anonymous):

I have no idea, I don't understand this.

OpenStudy (***[isuru]***):

Do you know wt is meant by sec value of an angle ? i.e - like sec (theta)

OpenStudy (anonymous):

Yes, secant. I just don't know how to setup.

OpenStudy (***[isuru]***):

First, u can't find the sec value of an angle without the calculator if u can't convert that value to a cos value of an angle i.e - \[\sec (\pi) = \frac{ 1 }{ \sec (\pi) }\] to find the value of sec (-7pi/3) without a cal u must write it as above .... Just don't think about the future steps. What will be ur angle equal to ?\[\sec(\frac{ -7\pi }{ 3 }) = \frac{ 1 }{ \cos [blank] }\]

OpenStudy (***[isuru]***):

\[\sec (\frac{ -7\pi }{ 3 }) = \frac{ 1 }{ \cos (\frac{ -7\pi }{ 3 } )} \] did ya get it ?